The diagonals of a rhombus are 10 centimeters

and 24 centimeters. Find the

perimeter of the rhombus.



1
Expert's answer
2021-02-02T05:15:33-0500

In any rhombus, the diagonals  bisect each other at right angles (90°). 

Assume "a" is the side of the rhombus, "d_1" and "d_2" are its diagonals

"d_1=10\ cm, d_2=24\ cm."

By the Pythagorean Theorem


"a^2=(\dfrac{d_1}{2})^2+(\dfrac{d_2}{2})^2"

"a=\sqrt{(\dfrac{10}{2})^2+(\dfrac{24}{2})^2}=13(cm)"

Find the perimeter of the perimeter


"P=4a=4(13 \ cm)=52\ cm"

The perimeter of the rhombus is 52 centimeters.